OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many solutions; Example:(2,1). x + y = 16 and x + y = -2 Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many solutions; Example:(2,1). x + y = 16 and x + y = -2 @Algebra

OpenStudy (asnaseer):

(1) x+y=16 (2) x+y=-2 This pair has no solution as x+y cannot have two different values.

OpenStudy (anonymous):

Ok show me this one step by step , Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many solutions; Example:(2,1). 3x - 2y = 9 and -3x + 2y = -9

OpenStudy (asnaseer):

(1) 3x-2y=9 (2) -3x+2y=-9 no multiply (2) by -1 to get: (3) 3x-2y=9 this is identical to (1). Therefore equations (1) and (2) are identical and you only have one equation with two unknowns. hence you cannot find a unique solution. hence there are infinitely many solutions.

OpenStudy (anonymous):

ok